OpenStudy (anonymous):

Sarah is the president of the drama club and wants to publish a newspaper. Expenses are $.90 for printing and mailing each copy, plus $600 total for research and writing. The price of the newsletter is $1.50 per copy. How many copies of the newsletter must the club sell to break even? a. To break even, the drama club must sell 1000 copies. b. To break even, the drama club must sell 100 copies. c. To break even, the drama club must sell 250 copies. d. To break even, the drama club must sell 444 copies.

OpenStudy (anonymous):

@bohotness help please

OpenStudy (anonymous):

i take help from people i know

OpenStudy (anonymous):

@bohotness fan me please it is eminem101

OpenStudy (anonymous):

@jigglypuff314

OpenStudy (anonymous):

@Lady.Liv1776 fan me please

jigglypuff314 (jigglypuff314):

Expense = 0.9(per copy) + 600 and Price = 1.5 (per copy) sooo 0.9x + 600 = 1.5x can you solve for x from there? :)

OpenStudy (anonymous):

no

OpenStudy (anonymous):

:(

jigglypuff314 (jigglypuff314):

here 0.9x + 600 = 1.5x try subtracting 0.9x from both sides -0.9x -0.9x ------------------ 600 = 0.6x do you see what I did there? :)

OpenStudy (anonymous):

kinda

jigglypuff314 (jigglypuff314):

then we can divide both sides by 0.6 and get x = 600 divided by 0.6 = ?

OpenStudy (anonymous):

is it 1000?

jigglypuff314 (jigglypuff314):

yep :)

OpenStudy (anonymous):

Julio has started his very own floral business. He sells centerpieces for $19.50 each. His fixed costs are $500 per month, and each centerpiece cost $8 to produce. Let n represent the number of centerpieces you sell. Write an equation to determine the cost, C. Write another equation to determine the revenue, R. a. 4072-03-01-07-00_files/i0200000.jpg b. 4072-03-01-07-00_files/i0200001.jpg c. 4072-03-01-07-00_files/i0200002.jpg d. 4072-03-01-07-00_files/i0200003.jpg

OpenStudy (anonymous):

does that work for the next question

jigglypuff314 (jigglypuff314):

please ask separate questions in a separate question post ^_^ and you can attach computer files with the "Attach File" button below :)

OpenStudy (anonymous):

i will close this one so i can make a new one ok